Consider a binomial experiment with n = 6 trials where the probability of success on a single trial is p = 0.35. (Round your answers to three decimal places.) (a) Find P(r = 0). (b) Find P(r ≥ 1) by using the complement rule.
Answer #1

Let X ~ Binomial ( n = 6, p = 0.35)

Probability mass function of X is,

P(X = r) = nCr pr (1 - p)n-r

We want to find the following probabilities,

a) P(r = 0)

= 6C0 * (0.35)0 * (1 - 0.35)6-0

= 1 * 1 * (0.65)6

= 0.075

=> P(r = 0) = 0.075

b) P(r ≥ 1) = 1 - P(r < 1) = 1 - P(r = 0) = 1 - 0.075 = 0.925

=> P(r ≥ 1) = 0.925